### v1

@@ 7 Conclusions and Future Work
For arbitrarily many players, we find socially optimal pure equilibria in some cases and bound the efficiency from below. For a better understanding of the existence and efficiency of NE for more than two players, we simulate fictitious play. To this end, we generalise fictitious play to infinite strategy spaces and describe some of the best responses of a player to the other players’ strategies. First, we corroborate all the theoretical predictions for the simulated cases. The most important factor for existence and efficiency of an equilibrium is the ratio of the largest to the second largest project function coefficients and of the largest to the second largest budgets. Therefore, to influence the projects and the players to some extent, one should influence the projects with the highest values and the players with the largest budgets.
@@
Consider directions for future work. Since the real value usually depends on the total contribution linearly up to a point, we should study piecewise project functions. More generally, we would like to extend our complete theoretical characterization of the existence of (pure) NE to more than two players and to various non-linear project functions. Next, we can try finding various NE that are not yet analysed analytically by approximately finding a best response and extending the simulations to more than two projects. Another numerical approach is discretizing the game and applying off-the-shelf software to find the equilibria of the discretization, such as Gambit  [^16] http://gambitproject.readthedocs.io/en/v15.1.1/intro.html# or Game Theory Explorer  [^17] http://banach.lse.ac.uk/. Since randomization can be deliberately undertaken or describe beliefs or uncertain behaviour (Osborne and Rubinstein, 1994, Section 3.2), we would like to find concrete mixed equilibria to be able to advise on playing them or at least predict the outcomes, like we do here for the pure equilibria.
@@ Appendix A Proofs
Aiming to bound the price of anarchy, we notice that since player $n$ can always obtain at least
@@
We shall now prove Theorem 10. To remain self-contained, before proving the theorem, we bring here the necessary definitions used by Maskin and Dasgupta (1986). Given player $i$ with the strategy set $A_{i}\subseteq\mathbb{R}^{m}$, define $A\stackrel{{\scriptstyle\rm\Delta}}{{=}}A_{1}\times\ldots\times A_{n}$.
@@
Proof. We show now that all the conditions of Theorem $5^{*}$ from Maskin and Dasgupta (1986) hold. First, the strategy set of player $i$ is simplex, and as such, it is non-empty, convex and compact. The utility function of player $i$ is discontinuous only at a threshold of one of the projects. These points belong to the set $A^{*}(i)$, defined in Formula(s) (5), if we take
@@
We now prove Theorem 8.
Proof. We reduce the following NP-hard problem (Martello and Toth, 1990, Section 1.3) to best-responding by, w.l.o.g., player $1$.

### v2

@@ 7 Conclusions and Future Work
For arbitrarily many players, we find socially optimal pure equilibria in some cases and bound the efficiency from below. We obtain further results on the existence and efficiency of NE for more than two players by simulating fictitious play. Still, the main contribution of the simulations resides in the presented methodology, generalising fictitious play and computing a best response. To this end, we generalise fictitious play to infinite strategy spaces and describe some of the best responses of a player to the other players’ strategies. First, we corroborate all the theoretical predictions for the simulated cases. The most important factor for existence and efficiency of an equilibrium is the ratio of the largest to the second largest project function coefficients and of the largest to the second largest budgets. Therefore, to influence the projects and the players to some extent, one should influence the projects with the highest values and the players with the largest budgets.
@@
Our main conceptual contribution is narrowing down a Nash equilibrium to a new solution concept, discovering that without suppression of whole budgets, Nash equilibria have no profitable cyclical deviations. That inspires thinking about other interactions and relevant solution concepts, which may coincide on important classes of games, motivating much future work.
Further directions for future work follow. Since the real value usually depends on the total contribution linearly up to a point, we should study piecewise project functions. Sometimes the value of the project grows in discrete steps; for example, in the times required to complete a product or to complete a certain part of an article. In those cases, replacing our linear model with a discrete value function would require us to check which properties still hold, and this more precise model would also inspire further realistic models in game theory. More generally, we would like to extend our complete theoretical characterization of the existence of (pure) NE to more than two players and to various non-linear project functions. Next, we can try finding various NE that are not yet analysed analytically by approximately finding a best response and extending the simulations to more than two projects. Our simulations possess extra range representativeness because of the multiplication invariance, and we have tried to consider the practically realistic ranges parameters. Still, the parameter range of the simulations should be extended based on the studied scenario. Another numerical approach is discretizing the game and applying off-the-shelf software to find the equilibria of the discretization, such as Gambit  [^14] http://gambitproject.readthedocs.io/en/v15.1.1/intro.html# or Game Theory Explorer  [^15] http://banach.lse.ac.uk/. Since randomization can be deliberately undertaken or describe beliefs or uncertain behaviour [40, Section 3.2], we would like to find concrete mixed equilibria to be able to advise on playing them or at least predict the outcomes, like we do here for the pure equilibria. Finally, we would like to vary the model, studying a relative threshold that depends on the median of the contributions, to dampen the influence of extreme contributors, or studying an absolute threshold, and imposing further constraints on the players. Converting important interactions, like investment policies of political campaigns, to our model opens a promising avenue, too.
@@ Appendix A Proofs
In order to bound the price of anarchy, we notice that since player $n$ can always obtain at least
@@
We shall now prove Theorem 9. To remain self-contained, before proving the theorem, we bring here the necessary definitions used by [32]. Given player $i$ with the strategy set $A_{i}\subseteq\mathbb{R}^{m}$, define $A\stackrel{{\scriptstyle\rm\Delta}}{{=}}A_{1}\times\ldots\times A_{n}$.
@@
Proof. We show now that all the conditions of Theorem $5^{*}$ from [32] hold. First, the strategy set of player $i$ is simplex, and as such, it is non-empty, convex and compact. The utility function of player $i$ is discontinuous only at a threshold of one of the projects. These points belong to the set $A^{*}(i)$, defined in Formula(s) (3), if we take
@@
Appendix B Simulations
The theory provides only sufficiency results for shared effort games with more than $2$ players. To better explore the existence and efficiency of NE in these games, we simulate a variation of fictitious play [45]; intuitively, we try fictitious play because of its well-known properties of converging to an NE in some finite cases (see Section 6). The main value of this section is in the methodology; the concrete simulations are best done based on the studied case at hand.
We first adapt the fictitious play [12] to our infinite game. Danskin [14] defines the best response to maximise the average utility against all the previous strategies of the other players, while we, as well as the original fictitious play, best respond to the averaged (cumulative) strategy of the others.
Aspiring to implement the adapted fictitious play, we then suggest an algorithm for finding a best response, if it exists, for the case of two projects. For more projects, we prove that even best responding is already NP-hard.
B.1 Infinite-Strategy Fictitious Play for Shared Effort Games
Definition 7.
Given a shared effort game with players $N$, budget-defined strategies $S^{i}=\left\{x^{i}=(x^{i}_{\omega})_{\omega\in\Omega}\in\mathbb{R}_{+}^{\left|\Omega\right|}\mid\sum_{\omega\in\Omega}{x^{i}_{\omega}}\leq B_{i}\right\}$ and utilities $u^{i}(x)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\sum_{\omega\in\Omega}{\phi^{i}_{\omega}(x_{\omega})}$, define an Infinite-Strategy Fictitious Play (ISFP) as the following set of sequences. Consider a (pure) strategy in this game at time $1$, i.e. $(x^{i}(1))^{i\in N}=((x^{i}(1)_{\omega})_{\omega\in\Omega})^{i\in N}$, define $X^{i}(1)\stackrel{{\scriptstyle\rm\Delta}}{{=}}\left\{x^{i}(1)\right\}$, and define recursively, for each $i\in N$ and $t\geq 0$, the set of the possible strategic histories at time $t+1$ to be all the possible combinations of the history with the best responses to the previous strategic histories:
In ISFP, all the plays have equal weights in the averaging. In the other extreme, a player just best-responds to the previous strategy profile of other players, thereby attributing the last play with the weight of $1$ and all the other plays with 00. In general, we define, for an $\alpha\in[0,\infty]$, an $\alpha-\text{ISFP}$ play as in Definition 7, but with the following formula instead [^16] For $\alpha=\infty$, we just obtain a constant sequence.
Here, the last play’s weight is $\frac{1}{\alpha t+1}$.
We do not know whether and when any convergence property can be proven for the generalised fictitious play in shared effort games. In simulations, our generalised fictitious play often converges to an NE, bolstering the importance of equilibrium efficiency, as stated by Roughgarden and Tardos [39, Chapter 17].
B.2 Best Response
Theorem 10.
Proof. We reduce the following NP-hard problem [31, Section 1.3] to best-responding by, w.l.o.g., player $1$.
@@ B.2 Best Response
In order to execute the infinite-strategy fictitious play, from now on we assume that the game has only two projects, $\Omega=\left\{\psi,\omega\right\}$. We would like to find a best response for a player $i\in N$, all the other players’ strategies $x^{-i}\in S^{-i}$ being fixed. From the weak monotonicity of the sharing functions $\phi^{i}_{\omega}$, we may assume w.l.o.g. that a best responding player contributes all her budget. Then, a strategy is uniquely determined by the contribution to project $\psi$ and we shall write $x^{i}$ for $x^{i}_{\psi}$, meaning that $x^{i}_{\omega}=B_{i}-x^{i}$.
Theorem 11.
B.3 The Simulation Method
[…]
B.4 Results and Conclusions